OpenStudy (anonymous):

Solve: 1700=300e^(r*11) for r

OpenStudy (anonymous):

I used the formula A=Pe^(rt) to form this equation

OpenStudy (anonymous):

1700/300=e^(r*11)

OpenStudy (anonymous):

1700/300 =e^(r*11) e^(r*11) =17/3 taking log both sides we have ln e^(r*11) = ln (17/3) 11r lne =ln17 -ln3 or r = (ln17 -ln3 )/11 as lne=1

OpenStudy (anonymous):

ok thank you very much. the key is to ln both sides and bring down the exponent then.

OpenStudy (anonymous):

yup welcome